The angular momentum of a flywheel having a rotational inertia 0.142kgm^2 decreases from 3.07 to 0.788kg-m^2/s in 1.53s. Find the average torque acting on it in this period.

Grade:12th pass

1 Answers

Khimraj
5 years ago
average torque will be
τ = I*\alpha = I*dw/dt
τ = 0.142*(3.07 – 0.788)/1.53 = 0.212 Nm
Hope it clears.